FEL Upgrade Project Weekly Brief – August 19-23, 2002




A technical requirements and preliminary design review was held this week (Aug. 22) on the laser microengineering experimental station that is being designed and built for the FEL Upgrade by the Aerospace Corporation as part of our joint contract with the AFRL. The review (which was held by teleconference to save travel time and costs) was judged to be successful.


On Aug. 19-20, Jefferson Lab held a review of the lab's Institutional plan for the new head of DOE's Office of Science, Ray Orbach. The OMS/AFRL funded FEL project was included in both the formal presentation and site tour.

The project technical and financial report for the month of July was completed and distributed to the DOE, ONR and AFRL contract monitors.

On Aug. 23, F. Dylla and G. Neil attended a briefing to the Deputy Secretary of the Navy (Livingstone) presented on behalf of Admiral Cohen (ONR) and Mathis (NAVSEA).

The dates for the next semiannual review of the FEL Upgrade project has been selected by the review committee: Nov. 20-21.

WBS 4 (Injector):

The SF6 tank was sent out for modifications. All vacuum screws required for gun assembly were submitted to the shop for venting. The support tube was made ready for E-beam welding of the support flange, the welding should be completed today. Drawings for the running resistor housing were signed off.

POISSON simulations of the two gun solenoids have been performed to study the effect of the increased aperture on the downstream solenoid. It is important to verify this because PARMELA input files have not been updated with such modification. So far, the simulations show a variation no larger than 2% in the peak magnetic field. A more detailed analysis is in progress.


Gun HVPS - The 3 stacks and their toroids are assembled in the HVPS Tank. The electronics is ready to mount in their holders on the side of the tank. The piping for the SF6 to the tank should be completed next week. A new Conditioning Resistor was made this week. It will be fired next week after it has dried more. Parts for the combined resistors and transmission line are being made by the Machine Shop.

WBS 6 (RF):

Quarter HVPS - The Crowbar unit is not sustaining its arc after it is triggered by a HV current surge. Discussions with EEV indicate the thyratron is worn out by usage (~10K hrs. of operation average). A spare thyratron is being installed and will be tested later today. The plumbing of LCW to the Circulator, RF Load, and HOM Filter continues.

Zone 4 - The half-height waveguide sections are being made in the Machine Shop with a delivery date of 8/30/02.

Zone 2 - The waveguide sweeps can be installed now. The cable pulls past this zone is complete.

Zone 1 - The waveguide can be reinstalled on the Quarter. The warm windows have been changed for higher RF power.

WBS 8 (Instrumentation):

The students are going back to school so the manpower crunch is on. Please say thanks to them if you see them - they all have worked their ***'s off this summer to make the FEL a success - Hat's Off!

A trip was made to Fermi Lab on Monday to review their new switched mode trim power supplies. They have developed a range of four quadrant power converter modules that cover the range of + 10 Amp, + 30 Amp, + 50 Amp, and + 100 Amp (2 50's in parallel). These are 7" high modules that can go 8 (10 Amp) across, the approach uses a single bulk supply at 150 Volts and FET switchers at 40KHz. We are getting a quote from a commercial design shop for adding a DSP/CAN interface to the control card.

BLMs - 1 of 6 of the BLMfront boards (F0151) has been partially populated and is undergoing some simple performance tests. Upon completion of the testing the board will be submitted back to have the next group of components installed. Everything is proceeding well. The work for ECO of the BLM rear cards (F0150) has been completed, however, the updated drawing packages have not been delivered as yet. DLPC - The software for the new pulse selection hardware is proceeding. In parallel, progress on the PCB layout for the CO-305 synchronous countdown (micropulse selection) carrier board continues as well. The new implementation of this device will provide the full capability of the con-optics into EPICS. MPS - The documentation for the 6U (vme) isolation board (GPIM64/05516-B-0009) has been collected and the components needed to assemble 4 to 6 of these cards are being procured.

The DSP/Pyroline camera project continues to make progress. This week we were able to output the dummy_clk_in from a timer (pretty much a noise free signal) and also able to get the dummy_pulse out of a GPIO port. Combined the two routines (a) generation of dummy_clk_in signal and (b) outputting a dummy_pulse waveform upon feeding the DSP with dummy_beam_sync. Testing has been good so far. Upon request from user (that's just temporary), A/D conversion can be achieved. We've been working on the part of inserting A/D with the rest of the 2 routines. If this starts to work then the timing issues need to be tended to next.

Design and fabrication of the gun High Voltage power supply interface is progressing. The new VME interface (Acromag card)channel assignments have been made are being programmed by Al. The PCB prototype for this system is nearing completion. The optical cavity wiring design is complete, Al is writing the software now. The interface chasssis design is complete but there is no manpower to start it yet.

The RS-485 cables have been terminated for the trim racks. Documentation is in the final stages for the 64 bit VME isolation daughter module for the MPS system. Power cables (535 MCM) were pulled from the 66 kW box supply to the to the wiggler, and 2 ga. cable has been pulled to sex/octapole magnet locations. Beam line has been installed in the dummy cryomodule section of the linac.

Further work has been completed on the Linux spot size code. The "background deletion" function now works with a two dimensional array, and the image capture and save functions have been debugged. Our current estimate on the performance of this program is 5-7 frames-per-second (fps), though we believe that can be improved through optimization of the capture code(the analysis functions are reasonably efficient). We will be getting in contact with Al Grippo soon to start implementing the EPICS part of this project. Code will be written to detect the tick marks, beam spot, and estimate the spot size, distribution, and position.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• The first GW underwent rudimentary magnet field tests for core flatness on the auxiliary magnet test stand. First data indicates core field is near the plus or minus 1 part in 10000 flatness over the good field region.

- The magnet runs cool, so the cooling scheme is OK.

• Contrary to the report last week, the remaining 6 cores of the second batch of 6 Magnets (for the UV line) had ambiguous gap dimension readings at Diamond Tool, the subcontract machining house. MEI is looking into it more closely.

Injector Dipoles (DU/DV)

• The measurements of the first GV continued. Ramp rate was slowed down to levels that are much less sensitive. The procedure for dialing in the right trim current for flat core field has become routine.

- The magnet runs cool, so the cooling scheme is OK.

- The first try at finding the main and trim current at various field levels corresponding to various injection energies was tried. The currents are almost linear and can be extrapolated to intermediate energies. Then a series of current trips at maximum current appeared to put the magnet in an entirely different remnant field configuration and we lost our repeatability. This looks like it will be our greatest challenge with this magnet.

- The first integrals were taken. The integrals are flat in the transverse direction to the 1 part in 10000 level. This attests to DULY Research's solution, using mu metal inserts in the field clamp's inner top surface and sliding shorts in the field clamp on either side of the beam tube.

- Integrals are about 1 % high at the present central field clamp positions. We will adjust the field clamps in to obtain the correct integral.

- Our next concentration is the value of the integrals on either side of the centerline of this wedge magnet.

• We are resolving our contract issues with Wang NMR. One GU and one GV will be air freighted today and the remaining five magnets will be shipped by truck.

Arc 180 Degree Dipoles (GY)

• GY Coil Potting fixture is almost ready.

• We are working out the contractual issues associated with winding a fifth coil to replace the potentially unreliable first coil.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Wang NMR received the GQ cores from Process Equipment Co.

• The GX cores are ready for their gap inspection.

• At Wang NMR, potting fixtures are almost ready for the GQ and GX coils.

• All GX and GQ coils are wound.


Trim Quad (QT)

• Milhous should be shipping two trim quads with the auxiliary cooling plates on the side coils. We are ready to test.

Sextupole (SF)

• DULY Research made substantial progress at finishing the Sextupole drawings. A second round of mark-ups were sent back to them with minor corrections.

• We are starting the order of the ceramic tube insulators that make the sextupole’s two water circuit per coil cooling scheme compact enough to fit within the magnet covers.

Octupole (OT)

• The Octupole design is about half complete at DULY Research while they finish the Sextupole.

Beam Line and Vacuum

• For the Arc Chamber job, Master Machine finally received the water jet cut materials. The effect of the delay is about three and a half weeks.

• Using the GW magnet, the test chambers were tested for their effect on core field and field flatness. The data is being analyzed.

• Using the lull in further fabrication of the injection line’s X Chamber (because of missing transition pieces), we tested the effect of the chamber cross section on field. The effect appears to be flat to within the resolution of the hall probe. The chamber welds may not require any machining to flatten the field. We will do one more position to verify the non-effect.

• We examined the inside welds of the X Chamber with a bore scope and found one region of insufficient penetration that required a small re-weld.

• The Jlab Shop continues fabrication of the four chambers for the optical chicane region.

• The layout of the drawing of the regions around the wiggler was signed off. The girders are now being detailed.

• The extraction "Y" chamber layout was signed off and the stand positions for the magnets were released to the alignment folks.

• The stands for sextupoles and trim quads are in fabrication and due in a week.

• As the vacuum crew came off CEBAF deployment, girder and vacuum assembly continued with the 5 F girders assembly and installation of the vacuum pipe between cryomodules.

• Low Conductivity Water piping installation continued in the FEL enclosure

• The alignment crew continued stand placement.

WBS 11 (Optics):

This week's mechanical work on the optical cavity was spent on reassembly of the optical cavity using a Si/Bronze nut with the sliver-sputtered leadscrew. We found that we still had high drag torque. In part this is due to the silver spalling off the leadscrew. We are preparing to try one of our options; the use of molybdenum disulfide (a JLab-approved UHV lubricant) sputtered onto a bare stainless steel leadscrew. This will be done next week. The HEPA filtered enclosure around the optical cavity test and assembly stands was completed.

After receiving feedback from the optics design group, we agreed that the optical transport turning cassette design was sound, and it is being detailed. As noted earlier, linear translators and LVDTs are already being procured. The collimator design is also going forward. Revisions to the optical cavity vacuum vessels (for the rotary feedthrough and the cooling water jacket) were signed this week.

Our optimizing the alignment of a photodiode that controls the modelocking of the ultrafast laser has improved the power stability and we are now getting consistent values of phase noise due to the PLL. Converted to rms timing jitter, values are no worse than 400 fs and with some judicious tradeoffs (e.g., output power), we can achieve values closer to 250 fs.

Other activities:

We provided input to Aerospace Corp. on their Systems Requirement Review Document. We participated in a review (through video teleconferencing) with them, and will address some action items that resulted. We spent some more time with potential vendors of the laser we'll use with the FEL external cavity (the pulse stacker).